OpenStudy (anonymous):

What is the solution set? Sq rt (3x+1) - sq rt (x-1) = 2

OpenStudy (anonymous):

\[\sqrt{3x+1} - \sqrt{x-1} = 2\]

OpenStudy (campbell_st):

square every term which gives 3x + 1 - x + 1 = 4 2x + 2 = 4 now solve for x